A sample of iron(III) oxide contains 6.7x1021 atoms of oxygen. What is its mass in grams
1
Expert's answer
2016-03-18T15:48:04-0400
Solution
N(O) = N/Na = 6.7*1021 * 6.02*1023 = 0.011 (mol)
N(Fe2O3) = N(O) / 3 = 0.011 / 3 = 0.0037 (mol)

Answer
N(Fe2O3) = 0.0037 mol